OpenStudy (anonymous):

The value of Peter's house has risen 10% a year for the last 3 years. If the value of Peter's home 3 years ago was $100,000, what is the value now?

OpenStudy (anonymous):

after one year value=100000+.01*100000=101,000 after two year value =101000+.01*101000=102,010 after 3 year value = 102010+.01*102010=$103030.1